# Learning the Basis: A Kolmogorov-Arnold Network Approach Embedding Green’s Function Priors

### v2

@@ IV Numerical Results and Discussion
TABLE I: Our Proposed PhyKAN Framework’s Reconstruction Performance.
Dataset Cube Sphere Cone Assembly Body
$B$-spline PhyKAN 0.0098 0.0089 0.0084 0.0092
Fourier PhyKAN 0.0123 0.0072 0.0101 0.0104
Rational PhyKAN 0.0115 0.0107 0.0119 0.0145
KA-GCN PhyKAN 0.0061 0.0075 0.0060 0.0064
KA-GAT PhyKAN 0.0107 0.0119 0.0079 0.0088
@@
To evaluate PhyKAN, MoM solutions are used as ground-truth for current reconstruction and RCS validation. The model inputs triangular mesh elements with EM descriptors and outputs induced currents. Tests are performed on four canonical PEC shapes [16] (cube, sphere, cone, assembly body) under a 1 GHz, unit-amplitude, vertical plane wave with $\theta,\phi\in[0^{\circ},180^{\circ}]$. Five variants assess the effect of KAN designs: three kernel activations ($B$-spline, Fourier, Rational) [18, 19, 20] and two graph KAN extensions (KA-GCN, KA-GAT) [21], used interchangeably in the local KAN blocks. Each is denoted [Kernel/Graph]-PhyKAN. DL-based CEM baselines either black-box the current mapping or only accelerate MoM without learnable bases; hence, we benchmark PhyKAN variants under a unified setting against MoM.
Figure 2 visualizes the surface reconstruction capability of five different variants of the proposed two-branch PhyKAN architecture. As observed, on the smooth geometry of the sphere, Fourier and KA-GCN PhyKAN accurately recover the global current distribution, while the $B$-spline and KA-GCN variants better capture sharp-edge behaviors on complex assemblies through localized and hierarchical aggregation. Quantitatively, Table I summarizes the mean-squared error (MSE) of each PhyKAN variant with respect to the MoM reference on the cube, sphere, cone, and assembly datasets. All variants show low reconstruction error against MoM, and the Fourier and KA-GCN configurations consistently achieve the highest accuracy across the canonical geometries. These results confirm that PhyKAN serves as a learnable, physics-adaptive basis, effectively overcoming the static rigidity of traditional RWG representations. This is actually attributed to its capability to tailor its current representation to the underlying geometry and field physics.
Figure 3: Bistatic RCS at $\phi=0^{\circ}$ for (a) the cube and (b) the cone objects, comparing different PhyKAN variants against the MoM reference.
TABLE II: Green’s Function Initialization (init.) Ablation for the Assembly Body Reconstruction.
Kernel MSE (no init.) MSE (Green init.)
KA-GCN PhyKAN 0.0123 0.0064
KA-GAT PhyKAN 0.0191 0.0088
To further assess the physical fidelity of the learned surface currents, the radar cross section (RCS) [24] is computed as a downstream validation in Fig. 3. It is shown that the predicted RCSs with all five considered PhyKAN variants agrees well with MoM references for both the cube and cone, even though this metric was never included as a training objective. This demonstrates that PhyKAN generalizes from near-field currents to physically consistent far-field responses, validating its physical consistency and confirming it avoids non-generalizable, black-box behavior.
In Table II, an ablation study, considering the assembly body, that directly validates the “Phy” component of the proposed PhyKAN framework is included. It can be seen that using Green’s function initialization significantly reduces reconstruction error. This confirms that embedding this physical prior is essential, resulting in strengthening both model accuracy and physical consistency. Overall, the results confirm that the proposed PhyKAN framework delivers geometry-adaptive and physically consistent surface current modeling, unifying data-driven flexibility with EM rigor.
@@ V Conclusion
This letter introduced PhyKAN, a physics-guided EFIE-derived framework that enables a learnable basis for EM modeling. By generalizing the RWG basis into an adaptive and interpretable representation, the proposed PhyKAN framework unifies diverse kernel and structural extensions within a physically consistent architecture. Future work will focus on scaling the learnable basis to large-scale and broadband scenarios.

### v3

@@ IV Numerical Results and Discussion
To evaluate PhyKAN, we assess its effectiveness, physical consistency, and adaptive behavior as a learnable basis. High-fidelity MoM solutions on a fine mesh ($\lambda/10$), the standard CEM reference [1], serve as ground truth for both surface-current reconstruction and Radar Cross Section (RCS) validation. The model takes triangular mesh elements with EM descriptors and predicts induced currents. Experiments cover four canonical PEC targets [16] (cube, sphere, cone, assembly body) and a complex SLICY object [25] under a 1 GHz, unit-amplitude, vertically polarized plane wave with $\theta,\phi\in[0^{\circ},180^{\circ}]$. DL-based CEM baselines either black-box the current mapping or only accelerate MoM without learnable bases; hence, we benchmark PhyKAN variants under a unified setting against MoM.
Figure 2: Surface current reconstruction on (a) the sphere and (b) the assembly body. Columns show the MoM reference, PhyKAN prediction, absolute error, KAN complexity maps, and the corresponding RCS comparison.
TABLE I: Relative MSE ($10^{-2}$) of PhyKAN and ablated variants.
Method Cube Sphere Cone Assembly Body
PhyKAN (no physics init) 3.42 2.95 3.88 6.73
PhyKAN (MLP Basis) 4.75 3.89 5.26 8.94
PhyKAN 1.21 0.94 1.28 1.84
TABLE II: Relative MSE ($10^{-2}$) of PhyKAN at different frequencies.
Frequency Cube Sphere Cone Assembly
1 GHz 1.21 0.94 1.28 1.84
2 GHz 1.46 1.07 1.53 2.21
5 GHz 1.89 1.33 1.98 2.67
Figure 2 shows the reconstruction results. Across both the sphere and the assembly body, PhyKAN closely matches the MoM current distributions, with absolute errors remaining small and localized. The KAN complexity maps quantify the average activation magnitude of all kernels in KAN blocks across the three edges of each element, yielding higher values near geometric discontinuities and thus highlighting PhyKAN’s adaptive allocation of $b_{n}(\cdot;\boldsymbol{u})$ compared with the fixed RWG basis. The far-field RCS results indicate that PhyKAN matches MoM patterns despite receiving no RCS supervision.
Table I reports the relative MSE for the ablation study. Removing the Green’s function initialization leads to higher errors across all geometries, while replacing the KAN basis with a plain MLP causes an even larger degradation. The full PhyKAN model achieves the best accuracy, demonstrating the complementary benefits of physics-based initialization and the learnable KAN basis. Table II summarizes PhyKAN’s accuracy when trained at 1, 2, and 5 GHz. The error increases only slightly with frequency due to finer high-frequency current oscillations, and remains stable across all targets, with a modest 1.3–1.6× rise from 1 to 5 GHz.
Fig. 3 evaluates performance under various mesh resolutions, using relative MSE with $\lambda/25$ MoM result as reference. In coarse-mesh regimes (e.g., $\lambda$, $\lambda/2.5$, $\lambda/5$), PhyKAN substantially reduces MSE on coarse meshes by compensating for geometric undersampling with its learnable basis. MoM overtakes only at very fine resolutions where computational cost dominates, making PhyKAN most valuable in coarse-mesh regimes.
Figure 3: Relative MSE ($10^{-2}$) of PhyKAN and MoM under different mesh resolutions, computed with respect to the $\lambda$/25 MoM reference.
@@ V Conclusion
This letter introduced PhyKAN, an EFIE-derived framework that replaces the fixed RWG basis with an adaptive learnable representation while embedding the Green’s function as a physics prior, forming a unified and physically consistent architecture for electromagnetic modeling. Future work will focus on scaling the learnable basis to large-scale and broadband scenarios.